OpenStudy (anonymous):

Solve. 5y+18+3y=14+8

OpenStudy (anonymous):

Get variables on one side and numbers on another! So 5y+3y=14+8-18. That'd leave you with 8y=4. Y=.5

OpenStudy (anonymous):

thank you